# Calibration of the strain amplitude recorded with DAS using a strainmeter array

### v1

@@ Acknowledgments
When using surface wave-data, which provide superior signal-to-noise ratio because of their large amplitudes (compared to body-wave signals) the limitations of the plane-wave approximation become significant. The surface waves are dispersive and do not propagate with a single phase-velocity and Love- and Rayleigh-waves are inseparably superimposed. Due to their propagation along the surface of the heterogeneous Earth’s crust, they have non-plane character and contain significant components, which do not propagate along the backazimuth (BAZ) great circle. These limitation become most obvious for BAZ being perpendicular to the azimuth of measured linear strain. The plane wave does not strain the material perpendicular to the direction of propagation and value of the cosine in Eq. (S6) necessarily vanishes. The recorded signals then entirely are composed by the non-plane components of the wave, not propagating along the BAZ great circle. The problem becomes already apparent for BAZ being close to perpendicular to strain azimuth. The seismometer recorded particle velocity contains as well these non-plane components, which are at variance with the expected polarization of the plane wave. Thus the scaling in Eq. (S6) is not valid for them. We discuss the consequences of this on the regression coefficients in section S4.2 ‘Regression with respect to seismometer data’. In section S4.3 ‘Comparison of seismometer and strainmeter data’ we discuss a direct comparison of strainmeter and seismometer data.
@@
Cases where BAZ is close to NE or NE have a great-circle propagation direction almost perpendicular to the azimuth of the ‘Anton Gang’ (NE). For the Vorstollen (NE), the same applies when BAZ is close to NE or NE. Most of these cases are illustrated in Fig. S32. The coefficients are significantly outside reasonable limits, with some even being negative. The scatter is more substantial for the ‘Anton Gang’, where we find several cases with BAZ almost perpendicular to the strain azimuth. For the ‘Vorstollen’ the study includes a case with BAZ near NE, but none with BAZ near NE. In consequence the overall scatter is less strong.
We also observe a systematic shift between coefficients obtained for regression with respect to strainmeter data compared to regression with respect to seismometer data. This might indicate that dominating signal propagates with a higher phase velocity than considered and that the value for the slowness in Eq. (S6) might be too large. It is, however, disputable, whether the entire difference of coefficients (a factor of 2 in some cases) in the ‘Vorstollen’ can be attributed to the wrong value of phase slowness, though there is no other obvious reason. The P-wave signals in Figs. S22, S23, S24, and S25 show an amplitude mismatch of a similar size. The mismatch varies along the P-wave-train and cannot be expressed by a single factor.
Hence, the issue is not limited to the regression coefficients. The DAS waveforms also show greater similarity to the strainmeter waveforms than to the seismometer waveforms. This is shown in Fig. S33, where the values of NCCs are plotted against backazimuth. The NNC measured by comparing DAS data with respect to seismometer data are generally smaller than for DAS with respect to the strainmeter data. For BAZ nearly perpendicular to the fiber optic cable azimuth, some of the signals are even anti-correlated (Fig. S33). This results from the change in the sign of the cosine in Eq. (S6).

### v2

@@ Acknowledgments
When using surface wave-data, which provide superior signal-to-noise ratio because of their large amplitudes (compared to body-wave signals) the limitations of the plane-wave approximation become significant. The surface waves are dispersive and do not propagate with a single phase-velocity and Love- and Rayleigh-waves are inseparably superimposed. Due to their propagation along the surface of the heterogeneous Earth’s crust, they have non-plane character and contain significant components, which do not propagate along the backazimuth (BAZ) great circle. These limitation become most obvious for BAZ being perpendicular to the azimuth of measured linear strain. The plane wave does not strain the material perpendicular to the direction of propagation and the value of the cosine in Eq. (7) necessarily vanishes. The recorded signals then entirely are composed by the non-plane components of the wave, not propagating along the BAZ great circle. The problem becomes already apparent for BAZ being close to perpendicular to strain azimuth. The seismometer recorded particle velocity contains as well these non-plane components, which are at variance with the expected polarization of the plane wave. Thus the scaling in Eq. (S6) is not valid for them. We discuss the consequences of this on the regression coefficients in section S4.2 ‘Regression with respect to seismometer data’. In section S4.3 ‘Comparison of seismometer and strainmeter data’ we discuss a direct comparison of strainmeter and seismometer data.
@@
Cases where BAZ is close to NE or NE have a great-circle propagation direction almost perpendicular to the azimuth of the ‘Anton Gang’ (NE). For the Vorstollen (NE), the same applies when BAZ is close to NE or NE. Most of these cases are illustrated in Fig. S34. The coefficients are significantly outside reasonable limits, with some even being negative. The scatter is more substantial for the ‘Anton Gang’, where we find several cases with BAZ almost perpendicular to the strain azimuth. For the ‘Vorstollen’ the study includes one case with BAZ near NE, but none with BAZ near NE. In consequence the overall scatter is less strong.
For data from the ‘Vorstollen’ we also observe a systematic shift between coefficients obtained for regression with respect to strainmeter data compared to regression with respect to seismometer data. The seismometer data appears to over-estimate the strain amplitude, such that it has to be downscaled by a factor of about 0.5 with respect to the coefficients obtained with the strainmeter data. The P-wave signals in Figs. S23 and S25 show an amplitude mismatch of a similar size. This might be a consequence of strain-strain coupling due to the local topography as discussed in section S2.5 ‘Distortion of the strain field due to local heterogeneity’. The East-West strain measured by the strainmeters as well as by the DAS fibers in the ‘Vorstollen’ is reduced compared to what is estimated from particle velocity (seismometer data). From tidal analysis Zürn et al. (2015, their table S2 in the supporting material) estimate a factor of 0.58 for NE and Emter and Zürn (1985, their figure 5) derive a factor of 0.67 by a 2D finite element analysis. Both are in the order of magnitude of the bias seen in Fig. S34 for the ‘Vorstollen’.
Apart from these amplitude related issues, the DAS waveforms also show greater similarity to the strainmeter waveforms than to the seismometer waveforms. This is shown in Fig. S35, where the values of NCCs are plotted against backazimuth. The NCC measured by comparing DAS data with respect to seismometer data are generally smaller than for DAS with respect to the strainmeter data. For BAZ nearly perpendicular to the fiber optic cable azimuth, some of the signals are even anti-correlated (Fig. S35). This results from the change in the sign of the cosine in Eq. (S6).